OpenStudy (anonymous):

The sum of Pete's and Sam's ages is 30. Five years ago, Pete was 3 times as old as Sam. How old is Sam? Let P = Pete's age, S = Sam's age, and P + S = 30. Which of the following equations would complete the system? A. P - 5 = 3S B. P - 5 = 3S - 5 C. P - 5 = 3S - 15

OpenStudy (broskishelleh):

P + S = 30 Five years ago, Peter's age : (P - 5) Sam's Age is: (S - 5) Now, (P - 5) = 3(S - 5) P - 5 = 3S - 15 P - 3S = -10

OpenStudy (anonymous):

Thank you! (: @Broskishelleh